# Entropy–Smooth Structures on Topological Manifolds

### v2

@@ Abstract.
We introduce an information–theoretic framework for smooth structures on topological manifolds, replacing coordinate charts with small-scale entropy data of local probability probes. A concise set of axioms identifies admissible coordinate functions and reconstructs a smooth atlas directly from the quadratic entropy response. We prove that this entropy–smooth structure is equivalent to the classical smooth structure, stable under perturbations, and compatible with products, submanifolds, immersions, and diffeomorphisms. This establishes smoothness as an information–theoretic phenomenon and forms the foundational layer of a broader program linking entropy, diffusion, and differential geometry.
@@ 1. Introduction
Differential geometry traditionally begins with smooth structure [lee2013smooth, hirsch1976differential]: a maximal atlas of compatible coordinate charts whose transition maps are $C^{\infty}$. This smooth atlas is imposed a priori and subsequently used to define tangent spaces, derivatives, differential forms, and geometric structures. In this paper we show that classical smooth structure can be characterized in terms of information–theoretic data: the small–scale entropy response of local information probes.
The starting point of this work is the information–theoretic program developed in our previous papers [sangha2025isoperimetric, sangha2025curvature], where we demonstrated that diffusion semigroups and entropy asymptotics encode Riemannian geometry. There, we established that a Riemannian manifold may be reconstructed from its diffusion behaviour: the metric norm arises from entropy expansion, Ricci curvature from convexity of entropy, and higher curvature data from higher–order information defects. Those results suggested that information flow is not merely a tool for analysing geometric structures but is sufficiently rich to faithfully recover them. This led naturally to the question motivating the present paper: can the smooth structure of a manifold be reconstructed from information alone, without any differential, metric, or geometric assumptions? Answering this question requires a new foundational framework. We introduce entropy–smooth manifolds: topological manifolds equipped with families of locally supported probability measures $\mu_{x,\varepsilon}$ that collapse to point masses as $\varepsilon\to 0$ and whose second–order entropy responses encode infinitesimal directions. Through a small set of natural axioms, we can identify coordinate functions by their entropy behaviour, assemble entropy coordinate charts, and generate an atlas whose compatibility follows from entropy calculus alone.
@@
The main results of the paper show that this information–theoretic definition of smoothness is equivalent to the classical one. Entropy coordinate charts are equivalent to classically defined $C^{\infty}$ coordinate charts; entropy transition maps are smooth; and the entropy–smooth functions coincide with classical smooth functions. We further establish that the quadratic entropy form determines the smooth structure uniquely, that products and submanifolds behave as expected, and that immersions, submersions, and diffeomorphisms admit precise entropy characterizations. Finally, entropy–smooth structures are stable under uniform weak convergence of probability measures.
Conceptually, entropy–smooth manifolds provide a new perspective on differentiable structure. Instead of postulating smooth charts, we detect smoothness through the response of information to local perturbations. This viewpoint is complementary to classical abstract frameworks such as Sikorski differential spaces or $C^{\infty}$–rings [sikorski1967differential, frolicher1986linear, moerdijk2003models] , which begin by specifying the algebra of smooth functions. Here, instead the smooth functions arise from the entropy structure itself, without assuming any differentiable structure in advance.
The results of this paper form the “zeroth–order” layer in our broader information–theoretic approach to geometry. Together with the author’s previous work on diffusion, curvature, and geometric analysis, they suggest a unified program in which classical and Riemannian geometry, smooth manifolds, and comparison theory all emerge from a single underlying principle: the structure of information flow at small scales. Future work will develop and connect these ideas with curvature, geometric flows, and the information–theoretic reconstruction of Riemannian geometry.
@@ 2. Basic definitions
For finite measures $\nu$ and $\mu$ on $X$, the Kullback–Leibler divergence [kullback1997information] is defined as
@@
[…]
$\nu_{x,\varepsilon,t}^{f}(A):=\int_{A}(1+tf)\,d\mu_{x,\varepsilon},\qquad A\subseteq X.$
Definition 2.2.
The entropy response is
$Ent_{x,\varepsilon}(t,f):=D(\nu_{x,\varepsilon,t}^{f}\|\mu_{x,\varepsilon}).$
The quadratic entropy response is
$I_{x,\varepsilon}(f):=\lim_{t\to 0}\frac{Ent_{x,\varepsilon}(t,f)-2Ent_{x,\varepsilon}(0,f)+Ent_{x,\varepsilon}(-t,f)}{t^{2}}$
whenever the limit exists in $0,\infty)$. The (small–scale) entropy coefficient is
$I_{x}(f):=\lim_{\varepsilon\to 0}I_{x,\varepsilon}(f).$
Note that the quadratic entropy response $I_{x,\varepsilon}$ is just the second derivative of the entropy response $Ent_{x,\varepsilon}(t,f)$ with respect to $t$ at 0:
$I_{x,\varepsilon}(f)=\left.\frac{d^{2}}{dt^{2}}\right|_{t=0}Ent_{x,\varepsilon}(t,f).$
@@
Definition 2.4.
For continuous functions $f,g:X\to\mathbb{R}$ and $\varepsilon>0$ we first define the $\varepsilon$–level joint entropy coefficient
$I_{x,\varepsilon}(f,g):=\lim_{t,s\to 0}\frac{D\bigl((1+tf+sg)\mu_{x,\varepsilon}\,\|\,\mu_{x,\varepsilon}\bigr)-D\bigl((1+tf)\mu_{x,\varepsilon}\,\|\,\mu_{x,\varepsilon}\bigr)-D\bigl((1+sg)\mu_{x,\varepsilon}\,\|\,\mu_{x,\varepsilon}\bigr)}{ts},$
whenever this limit exists. We require that, for entropy–smooth functions $f,g$, these limits exist for each $\varepsilon>0$ and that the convergence in $(t,s)$ is uniform as $\varepsilon\to 0$.
The joint entropy coefficient is then defined by
@@
whenever this limit exists in $\mathbb{R}$.
@@ 3. Axioms for Entropy–Smoothness
which is well–defined by viewing $A$ as an index set. Condition (3) implies that $\Phi$ is a topological embedding, and (1)–(2) imply that any component of $\Phi$ is obtained from the coordinate map $z\mapsto z$ by smooth composition. One checks that the image $\Phi(\Omega)$ is an embedded submanifold of $\mathbb{R}^{A}$ with the unique smooth structure for which all coordinates $f\in A$ are smooth, and that this structure coincides with the standard smooth structure on $\Omega\subset\mathbb{R}^{n}$ via the inverse of $\Phi$. In particular, every $f\in A$ is classically smooth on $\Omega$, so $A\subset C^{\infty}(\Omega)$. For details, we refer to [sikorski1967differential] or [Nestruev03]. ∎
@@
Let $\varphi:\widetilde{U}\to\mathbb{R}$ be bounded and continuous, and put $\widehat{\varphi}:=\varphi\circ F:W\to\mathbb{R}$. For fixed $z=F(x)$ and $\varepsilon>0$ we have
$\nu^{\,\varphi}_{z,\varepsilon,t}=(1+t\varphi)\,\widetilde{\mu}_{z,\varepsilon}=(1+t\varphi)\,F_{\#}\mu_{x,\varepsilon}=F_{\#}\bigl((1+t\widehat{\varphi})\mu_{x,\varepsilon}\bigr),$
so by invariance of Kullback–Leibler divergence under pushforward by a measurable bijection,
@@
Passing first to the quadratic entropy response and then to the small–scale limit gives
@@
whenever either side is finite. In particular, $\varphi$ is entropy–smooth on $\widetilde{U}$ if and only if $\widehat{\varphi}=\varphi\circ F$ is entropy–smooth on $W$.
@@ 3.1. Discussion of the Axioms
In summary, entropy–smoothness captures smooth structure on a manifold without assuming any differentiability on $X$ itself. The only use of classical smoothness occurs in Axiom 4 on regularity, where the smoothness of functions $h:\mathbb{R}^{k}\to\mathbb{R}$ is invoked to enforce the correct algebraic closure that characterizes $C^{\infty}$ structures. This reliance on Euclidean smoothness seems unavoidable and is shared by all alternative frameworks for abstract smooth structures, such as differential spaces [sikorski1967differential], diffeological spaces [IglesiasZemmour13, Souriau80], $C^{\infty}$–rings [moerdijk2003models, Kock06, JoyceCinfRings], convenient calculus [KrieglMichor97], synthetic differential geometry [Kock06], Frölicher spaces, and related approaches [Nestruev03]. All known abstract smooth-structure frameworks ultimately derive their notion of smoothness from smoothness on Euclidean spaces. Indeed, smoothness on $\mathbb{R}^{n}$ is taken as primary, and smoothness on arbitrary spaces is defined relative to it.
@@ 3.2. Correspondence with diffusion semigroups
If $(P_{t})_{t>0}$ is a diffusion semigroup on $X$, put $\mu_{x,\varepsilon}:=P_{\varepsilon}(x,\cdot)$. Here $P_{\varepsilon}(x,\cdot)$ denotes the diffusion kernel at time $\varepsilon$, i.e. a probability measure on $X$ depending on the starting point $x$. Then $\mu_{x,\varepsilon}$ describes the law of the diffusion started at $x$ after time $\varepsilon$. Our axioms express smoothness as the stability of relative entropy under infinitesimal perturbations of these transition probabilities. See [bakry2014analysis] for foundational material.
@@ 3.3. Correspondence with classical mollifiers
On a classical smooth manifold, choose a coordinate chart and a smooth compactly supported mollifier kernel $\eta_{\varepsilon}$ [lee2013smooth]. Push it forward to $X$ using the chart and glue globally using a partition of unity. The resulting measures $\mu_{x,\varepsilon}$ approximate $\delta_{x}$ as $\varepsilon\to 0$. Thus the entropy–smooth axioms correspond classically to smooth approximate identities.
@@ 4. Examples
Fix $x\in\mathbb{R}$. Let $dy$ denote Lebesgue measure on $\mathbb{R}$, and let the local information probes be the centred Gaussian measures
@@
For a bounded continuous function $f:\mathbb{R}\to\mathbb{R}$, the perturbed measure is $\nu^{\,f}_{x,\varepsilon,t}=(1+tf)\mu_{x,\varepsilon}$, and a direct Taylor expansion gives
$D\bigl((1+tf)\mu_{x,\varepsilon}\,\|\,\mu_{x,\varepsilon}\bigr)=\frac{t^{2}}{2}\int_{\mathbb{R}}f(y)^{2}\,\mu_{x,\varepsilon}(dy)+o(t^{2})\quad\text{as }t\to 0.$
Hence
$I_{x,\varepsilon}(f)=\int_{\mathbb{R}}f(y)^{2}\,\mu_{x,\varepsilon}(dy),\qquad I_{x}(f)=\lim_{\varepsilon\to 0}\int_{\mathbb{R}}f(y)^{2}\,\mu_{x,\varepsilon}(dy)=f(x)^{2},$
since $\mu_{x,\varepsilon}\Rightarrow\delta_{x}$ weakly.
In particular, for any smooth coordinate chart $F:U\to\mathbb{R}$ on an open interval $U\subset\mathbb{R}$, we obtain
$I_{x}(F)=F(x)^{2}.$
Given any point $x_{0}\in\mathbb{R}$, we may choose a local coordinate $F_{x_{0}}(y):=F(y)-F(x_{0})+c$ with $c\neq 0$, so that $F_{x_{0}}(x_{0})=c$ and hence $I_{x_{0}}(F_{x_{0}})=c^{2}>0$. Thus, on $\mathbb{R}$ the quadratic entropy coefficient recovers the usual one–dimensional smooth structure, and entropy coordinate charts can be chosen so that the $1\times 1$ information Gram matrix is strictly positive.

### v3

# Entropy-Smooth Structures on Topological Manifolds

@@ Abstract.
We present an information–theoretic characterization of smooth structures on topological manifolds. Instead of postulating a smooth atlas a priori, we consider families of local probability measures whose small–scale entropy response encodes infinitesimal directional information. From a concise set of axioms governing this quadratic entropy response, we identify admissible coordinate functions, construct entropy coordinate charts, and assemble a smooth atlas whose transition maps are smooth. We prove that the resulting entropy–smooth structure is equivalent to the classical smooth structure: entropy–smooth functions coincide with smooth functions, entropy coordinate charts are smooth charts, and the induced category is equivalent to the category of smooth manifolds. The framework is stable under perturbations and compatible with products, submanifolds, immersions, submersions, and diffeomorphisms. This establishes smooth structure as an information–theoretic invariant under explicitly stated axioms, and forms the zeroth–order layer of a broader program connecting entropy, diffusion, and differential geometry.
@@ 1. Introduction
Differential geometry traditionally begins by postulating smooth structure: a maximal atlas of coordinate charts whose transition maps are smooth. This smooth atlas is imposed a priori and subsequently used to define tangent spaces, derivatives, differential forms, and geometric structures. From a foundational point of view, it is natural to ask whether smooth structure itself admits an intrinsic characterization, independent of coordinates and differential calculus.
In this paper we address this question from an information–theoretic perspective. Rather than attempting to derive smoothness from minimal or purely topological assumptions, we introduce an axiomatic framework in which smooth structure is characterized by the small–scale behaviour of entropy associated with local probability probes. The central object is a quadratic entropy response that measures local variation of functions under infinitesimal perturbations of these probes. Under a precise and explicit set of axioms, this entropy response identifies admissible coordinate functions, yields entropy coordinate charts, and determines a smooth atlas.
The purpose of this work is not to claim that arbitrary topological manifolds admit smooth structures detectable from entropy alone. Instead, we show that whenever a topological manifold is equipped with families of local information probes satisfying natural locality, stability, and non–degeneracy axioms, the resulting entropy response encodes exactly the same information as a classical smooth structure. In this sense, entropy–smoothness provides an information–theoretic characterization of smooth manifolds. The role of the axioms is to make explicit the minimal information–theoretic content required to recover smoothness, rather than to conceal differentiability assumptions in analytic or geometric form.
Conceptually, the entropy–smooth framework is complementary to abstract approaches to smooth structure such as differential spaces, smooth function algebras, and related categorical formulations. Those frameworks begin by postulating an algebra of smooth functions closed under suitable operations. Here, by contrast, the class of smooth functions is selected by the behaviour of entropy at small scales, using only probabilistic and information–theoretic primitives. Closure under smooth functional composition appears explicitly as an axiom, reflecting the fact that smoothness on Euclidean space remains the universal reference point for any abstract smooth structure.
@@
The main results of the paper show that this information–theoretic notion of smoothness is equivalent to the classical one. Entropy coordinate charts are smooth diffeomorphisms onto their images, entropy transition maps are smooth, and entropy–smooth functions coincide with classical smooth functions. As a consequence, the category of entropy–smooth manifolds is equivalent to the category of smooth manifolds. The framework is stable under perturbations of the information probes and is compatible with products, submanifolds, immersions, submersions, and diffeomorphisms.
This work forms the zeroth–order layer of a broader information–theoretic approach to geometry developed in the author’s previous papers, where diffusion semigroups and entropy asymptotics were shown to encode Riemannian metrics and curvature. Here we isolate smooth structure itself, prior to any metric or geometric data, and show that it too can be captured by entropy at the smallest scales. Subsequent work will connect the entropy–smooth framework developed here to the information–theoretic reconstruction of Riemannian geometry and curvature, and to the development of entropy–based differential operators.
@@ 2. Basic definitions
For finite measures $\nu$ and $\mu$ on $X$, the Kullback–Leibler divergence [6] is defined as
@@
Let $f:X\to\mathbb{R}$ be bounded and continuous. The key lesson of the audit is that multiplicative perturbations of a fixed probe, $(1+tf)\mu_{x,\varepsilon}$, only measure values of $f$ at small scale and therefore cannot encode $n$ independent infinitesimal directions when $n>1$. To obtain genuinely local directional information from KL divergence we instead use exponential tilts of increments.
Increment normalization. For $x\in X$, $\varepsilon>0$, and a bounded function $f$, define the normalized increment
$\delta_{x,\varepsilon}f(y):=\frac{f(y)-f(x)}{\sqrt{\varepsilon}}.$
Definition 2.2 (Entropic tilt).
Fix $x\in X$ and $\varepsilon>0$. For a bounded continuous $f:X\to\mathbb{R}$ and $t\in\mathbb{R}$ define the (centered, normalized) exponential tilt of $\mu_{x,\varepsilon}$ by
$\frac{d\mu^{t,f}_{x,\varepsilon}}{d\mu_{x,\varepsilon}}(y):=\frac{\exp\bigl(t\,\delta_{x,\varepsilon}f(y)\bigr)}{Z_{x,\varepsilon}(t,f)},\qquad Z_{x,\varepsilon}(t,f):=\int_{X}\exp\bigl(t\,\delta_{x,\varepsilon}f\bigr)\,d\mu_{x,\varepsilon}.$
Definition 2.3 (Quadratic KL response).
For bounded continuous $f:X\to\mathbb{R}$, define the KL response
$\mathrm{Ent}_{x,\varepsilon}(t,f):=D\bigl(\mu^{t,f}_{x,\varepsilon}\,\|\,\mu_{x,\varepsilon}\bigr).$
The (increment-based) quadratic entropy response is
$I_{x,\varepsilon}(f):=\left.\frac{d^{2}}{dt^{2}}\right|_{t=0}\mathrm{Ent}_{x,\varepsilon}(t,f)\in[0,\infty]$
whenever the derivative exists. The (small–scale) entropy coefficient is
$I_{x}(f):=\lim_{\varepsilon\to 0}I_{x,\varepsilon}(f),$
when the limit exists and is finite.
Remark 2.4 (What $I_{x,\varepsilon}$ measures).
A standard cumulant computation shows
$I_{x,\varepsilon}(f)=\mathrm{Var}_{\mu_{x,\varepsilon}}\bigl(\delta_{x,\varepsilon}f\bigr)=\frac{1}{\varepsilon}\int_{X}\bigl(f(y)-f(x)\bigr)^{2}\,d\mu_{x,\varepsilon}(y)-\Bigl(\frac{1}{\sqrt{\varepsilon}}\int_{X}\bigl(f(y)-f(x)\bigr)\,d\mu_{x,\varepsilon}(y)\Bigr)^{2},$
so $I_{x,\varepsilon}$ measures local variation of $f$ at scale $\varepsilon$. Unlike the multiplicative model, this does not collapse to $f(x)^{2}$.
@@
Definition 2.6 (Joint entropy coefficient).
For bounded continuous $f,g:X\to\mathbb{R}$ and $\varepsilon>0$ define the $\varepsilon$–level joint entropy coefficient by
$I_{x,\varepsilon}(f,g):=\left.\frac{\partial^{2}}{\partial t\,\partial s}\right|_{(t,s)=(0,0)}D\bigl(\mu^{tf+sg,\,(f,g)}_{x,\varepsilon}\,\|\,\mu_{x,\varepsilon}\bigr),$
where the two–feature tilt is
$\frac{d\mu^{tf+sg,\,(f,g)}_{x,\varepsilon}}{d\mu_{x,\varepsilon}}(y):=\frac{\exp\bigl(t\,\delta_{x,\varepsilon}f(y)+s\,\delta_{x,\varepsilon}g(y)\bigr)}{\int_{X}\exp\bigl(t\,\delta_{x,\varepsilon}f+s\,\delta_{x,\varepsilon}g\bigr)\,d\mu_{x,\varepsilon}}.$
Whenever this exists, one has the covariance identity
$I_{x,\varepsilon}(f,g)$ $=\mathrm{Cov}_{\mu_{x,\varepsilon}}\bigl(\delta_{x,\varepsilon}f,\,\delta_{x,\varepsilon}g\bigr)=\frac{1}{\varepsilon}\int_{X}\bigl(f(y)-f(x)\bigr)\bigl(g(y)-g(x)\bigr)\,d\mu_{x,\varepsilon}(y)$
- $\Bigl(\frac{1}{\sqrt{\varepsilon}}\int_{X}\bigl(f(y)-f(x)\bigr)\,d\mu_{x,\varepsilon}(y)\Bigr)\Bigl(\frac{1}{\sqrt{\varepsilon}}\int_{X}\bigl(g(y)-g(x)\bigr)\,d\mu_{x,\varepsilon}(y)\Bigr).$
For entropy–smooth $f,g$ we define
@@
when the limit exists in $\mathbb{R}$.
@@ 3. Axioms for Entropy–Smoothness
which is well–defined by viewing $A$ as an index set. Condition (3) implies that $\Phi$ is a topological embedding, and (1)–(2) imply that any component of $\Phi$ is obtained from the coordinate map $z\mapsto z$ by smooth composition. One checks that the image $\Phi(\Omega)$ is an embedded submanifold of $\mathbb{R}^{A}$ with the unique smooth structure for which all coordinates $f\in A$ are smooth, and that this structure coincides with the standard smooth structure on $\Omega\subset\mathbb{R}^{n}$ via the inverse of $\Phi$. In particular, every $f\in A$ is classically smooth on $\Omega$, so $A\subset C^{\infty}(\Omega)$. For details, we refer to [10] or [9]. ∎
@@
Let $\varphi:\widetilde{U}\to\mathbb{R}$ be bounded and continuous, and put $\widehat{\varphi}:=\varphi\circ F:W\to\mathbb{R}$. Fix $z\in\widetilde{U}$ and write $x:=F^{-1}(z)\in W$.
We claim that the quadratic entropy coefficients on $\widetilde{U}$ computed using the pushed–forward probes $\widetilde{\mu}_{z,\varepsilon}=F_{\#}\mu_{x,\varepsilon}$ satisfy
@@
whenever either side is finite.
To see this, consider the entropic tilt on $\widetilde{U}$:
$\frac{d\widetilde{\mu}^{t,\varphi}_{z,\varepsilon}}{d\widetilde{\mu}_{z,\varepsilon}}(u):=\frac{\exp\bigl(t\,\delta_{z,\varepsilon}\varphi(u)\bigr)}{\widetilde{Z}_{z,\varepsilon}(t,\varphi)},\qquad\delta_{z,\varepsilon}\varphi(u)=\frac{\varphi(u)-\varphi(z)}{\sqrt{\varepsilon}}.$
For $u=F(y)$ with $y\in W$ we have $\varphi(u)=\widehat{\varphi}(y)$ and $\varphi(z)=\widehat{\varphi}(x)$, hence
$\delta_{z,\varepsilon}\varphi\bigl(F(y)\bigr)=\delta_{x,\varepsilon}\widehat{\varphi}(y).$
It follows that the tilted measure is the pushforward of the tilted measure on $W$:
$\widetilde{\mu}^{t,\varphi}_{z,\varepsilon}=F_{\#}\mu^{t,\widehat{\varphi}}_{x,\varepsilon}$,
@@
[…]
Taking the second derivative at $t=0$ and then the limit $\varepsilon\to 0$ yields (1). In particular, $\varphi$ is entropy–smooth on $\widetilde{U}$ if and only if $\widehat{\varphi}$ is entropy–smooth on $W$.
@@
Thus $\Phi:\widetilde{U}\to\widetilde{V}$ has classically $C^{\infty}$ components.
@@ 3.1. Discussion of the Axioms
In summary, entropy–smoothness captures smooth structure on a manifold without assuming any differentiability on $X$ itself. The only use of classical smoothness occurs in Axiom 4 on regularity, where the smoothness of functions $h:\mathbb{R}^{k}\to\mathbb{R}$ is invoked to enforce the correct algebraic closure that characterizes $C^{\infty}$ structures. This reliance on Euclidean smoothness seems unavoidable and is shared by all alternative frameworks for abstract smooth structures, such as differential spaces [10], diffeological spaces [2, 11], $C^{\infty}$–rings [8, 4, 3], convenient calculus [5], synthetic differential geometry [4], Frölicher spaces, and related approaches [9]. All known abstract smooth-structure frameworks ultimately derive their notion of smoothness from smoothness on Euclidean spaces. Indeed, smoothness on $\mathbb{R}^{n}$ is taken as primary, and smoothness on arbitrary spaces is defined relative to it.
@@ 3.2. Correspondence with diffusion semigroups
If $(P_{t})_{t>0}$ is a diffusion semigroup on $X$, put $\mu_{x,\varepsilon}:=P_{\varepsilon}(x,\cdot)$. Here $P_{\varepsilon}(x,\cdot)$ denotes the diffusion kernel at time $\varepsilon$, i.e. a probability measure on $X$ depending on the starting point $x$. Then $\mu_{x,\varepsilon}$ describes the law of the diffusion started at $x$ after time $\varepsilon$. Our axioms express smoothness as the stability of relative entropy under infinitesimal perturbations of these transition probabilities. See [1] for foundational material.
@@ 3.3. Correspondence with classical mollifiers
On a classical smooth manifold, choose a coordinate chart and a smooth compactly supported mollifier kernel $\eta_{\varepsilon}$ [7]. Push it forward to $X$ using the chart and glue globally using a partition of unity. The resulting measures $\mu_{x,\varepsilon}$ approximate $\delta_{x}$ as $\varepsilon\to 0$. Thus the entropy–smooth axioms correspond classically to smooth approximate identities.
@@ 4. Examples
Fix $x\in\mathbb{R}$. Let $dy$ denote Lebesgue measure on $\mathbb{R}$, and take the local information probes to be the centred Gaussian measures
@@
For a bounded continuous function $f$, our increment normalization is $\delta_{x,\varepsilon}f(y)=(f(y)-f(x))/\sqrt{\varepsilon}$. A Taylor expansion of the cumulant $\log Z_{x,\varepsilon}(t,f)$ gives
$I_{x,\varepsilon}(f)=\mathrm{Var}_{\mu_{x,\varepsilon}}\bigl(\delta_{x,\varepsilon}f\bigr)=\frac{1}{\varepsilon}\int_{\mathbb{R}}(f(y)-f(x))^{2}\,\mu_{x,\varepsilon}(dy).$
If $f$ is classically $C^{1}$ near $x$, then $f(y)-f(x)=f^{\prime}(x)(y-x)+o(|y-x|)$, and since $\mathrm{Var}(y-x)=\varepsilon$ under $\mu_{x,\varepsilon}$ we obtain
$I_{x}(f)=\lim_{\varepsilon\to 0}I_{x,\varepsilon}(f)=\bigl(f^{\prime}(x)\bigr)^{2}.$
In particular, the identity coordinate $F(y)=y$ satisfies $I_{x}(F)=1$ for all $x$, so the one–dimensional coordinate non–degeneracy axiom holds and the resulting entropy coordinate charts recover the usual smooth structure on $\mathbb{R}$.